# A Representation Transformation of Parametric Feynman Integrals

### v2

@@ Abstract
A transformation on homogeneous polynomials is proposed, which is further applied to parametric Feynman integrals. The two representations related through this transformation are dual to each other. And it naturally leads to dualities of Landau equations and linear integral relations between the two representations. For integrals with momentum-space correspondences, the dual representation is equivalent to the Baikov representation.
@@ I Introduction
Feynman integrals are a crucial element of perturbative quantum field theory. There are several representations of Feynman integrals, with the most well-known being the Feynman-parameter representation [1, 2, 3, 4]. Variants of the Feynman-parameter representation include the Lee-Pomeransky representation [5, 6] and the version used by the author in refs. [7, 8, 9, 10]. The parametric representation is powerful in the sense that it allows for both direct integration [11, 12, 13, 14, 15] and integral reduction [16, 6, 17, 7, 8, 9, 18, 19, 20], and Landau equations [21, 22, 23] can easily be formulated within this framework (see e.g. ref. [24]). Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], coaction [38, 39, 40, 41, 42], etc.
Another widely used representation is the Baikov representation [43, 44, 45, 46, 47] and its variants [48, 49, 50]. Although it is less suited for direct integration, the Baikov representation has the advantage that implementing cuts is straightforward [48, 51, 52], and it is more compatible with twisted cohomology theory [53, 54, 55, 56]. The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc..
While it is evident that these two representations should be equivalent, since both are derived from the momentum representation, it was not clear how to directly derive one representation from the other. This problem is addressed in this article. It is demonstrated that these two representations are related by a simple transformation. A transformation on homogeneous polynomials is introduced in section II. The transformation is then applied to parametric Feynman integrals in section III. It is proven that the two representations related through this transformation are dual to each other. The equivalence between the dual representation and the Baikov representation is proven in section IV.
@@ II Polynomial transformation
We assume that the degree of $\mathcal{F}$ is $L+1$. Then obviously
@@
Taking the derivative of both sides with respect to $u_{i}$, we get
@@
Generally speaking, the transformation defined by eq. II.1a may not be reversible. Even if it is, the obtained function $\mathcal{G}(u)$ may not be a rational function. However, for the $\mathcal{F}$ polynomial of a loop integral with a complete set of propagators, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
@@ III Representation transformation
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})\\ =&\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
@@
$w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v-i0^{+})^{\lambda+1}}e^{ivx}~{}.$ (III.2)
@@
$w_{-1}(x)=$ $2\pi\delta(x)~{},$
$w_{\lambda}(x)=$ $2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
And we have
$\frac{1}{x^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(x)e^{ivx}~{}.$ (III.3)
Obviously we can write $\frac{1}{\Gamma(\lambda+1)}\int_{0}^{\infty}\mathrm{d}xx^{\lambda}$ as $\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}xw_{\lambda}(x)$. By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$], it is easy to get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
where the last step can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
@@
$\Lambda_{i,j(n+1)}=$ $0,~{}j\neq n+1~{},$ (III.6a)
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
with $n_{0}$ a constant. As can be seen in appendix B, normal loop integrals do satisfy these conditions. The first condition is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$. For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eq. III.6, it can be proven that (see appendix A)
$\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}^{\prime}}~{},$ (III.7)
where $\lambda_{0}^{\prime}\equiv\lambda_{0}+\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a constant. Thus we get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}\\ \equiv&(2i\pi)^{-n-1}\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(\lambda_{0}^{\prime})}\\ &\times I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$.
@@
$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{iux}$
Table 1: The duality between the two representations.
@@ IV The Baikov representation
We consider a loop integral with a complete set of propagators
@@
where $l$ are the loop momenta and $p$ are the external momenta. For simplicity, I assume that all the propagators are linearly independent and the Gram determinant of the external momenta $p$ does not vanish. For future convenience, we collectively denote $l$ and $p$ by $q$. Scalar integrals can be parametrized by integrals of the form [44, 45]
$J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},$ (IV.2)
where $E$ is the number of the external momenta $p$. Here we have omitted some irrelevant prefactors for brevity. The polynomial $P(v)$ is proportional to the Gram determinant of the momenta $q$. That is
@@
In this section, we will prove that $J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})$ is equivalent to the integral $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ in eq. III.8 up to some constants.
Taking $\mathcal{E}^{(1)}=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
@@
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{-\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
×$\times\left[\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n}^{\prime}+1)}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]$
$=$ $2\Gamma(\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\frac{-4i}{\Gamma(1-\lambda_{0}^{\prime})}\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)$
×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
where $\mathcal{G}_{+}\equiv\left.\mathcal{G}\right|_{u_{n+1}=1}$ and $\mathcal{G}_{-}\equiv\left.\mathcal{G}\right|_{u_{n+1}=-1}$. Due to the positivity of $|u_{n+1}|$, the integral recieves contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$. The second step is carried out by reversing the path of integration for all $x_{i}$ for the first term in the squared bracket. And in the last step, we have closed the contour around the branch cut. Notice that we use such a convention that $\mathrm{Im}\{\mathcal{F}\}<0$. Thus $\mathrm{Im}\{\mathcal{G}\}>0$ when $\mathrm{Re}\{\mathcal{G}\}<0$. Now the region of integration is constrained by the theta function $\theta\left(-\mathcal{G}_{-}\right)$. Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=\lambda_{0}+\frac{1}{2}(L+E+1)$. Comparing the above integral with the Baikov integral in eq. IV.2, we see that these two integrals are equivalent if
$P(v_{1},v_{2},\dots,v_{n})=-\mathcal{G}_{-}=-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.$ (IV.4)
@@
$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6)
$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.7)
$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.8)
$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.9)
@@
$\begin{split}\mathcal{G}=&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.10)
@@
$P=$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
$=$ $\det\left(l_{i}\cdot l_{j}-\sum_{mn}g_{mn}l_{i}\cdot p_{m}l_{j}\cdot p_{n}\right)~{}.$
It is easy to get
@@
$P=\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.$ (IV.11)
Comparing eq. IV.10 with eq. IV.11, we get eq. IV.4.

### v3

@@ Abstract
A transformation on homogeneous polynomials is proposed, which is further applied to parametric Feynman integrals. The two representations related through this transformation are dual to each other. It naturally leads to dualities of Landau equations and linear integral relations between the two representations. For integrals with momentum-space correspondences, the dual representation is equivalent to the Baikov representation.
@@ I Introduction
Feynman integrals are a crucial element of perturbative quantum field theory. There are several representations of Feynman integrals, with the most well-known being the Feynman-parameter representation [1, 2, 3, 4]. Variants of the Feynman-parameter representation include the Lee-Pomeransky representation [5, 6] and the version used by the author in refs. [7, 8, 9, 10]. The parametric representation is powerful in the sense that it allows for both direct integration [11, 12, 13, 14, 15] and integral reduction [16, 6, 17, 7, 8, 9, 18, 19, 20], and Landau equations [21, 22, 23] can easily be formulated within this framework (see e.g. ref. [24]). Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], the coaction [38, 39, 40, 41, 42], etc.
Another widely used representation is the Baikov representation [43, 44, 45, 46, 47] and its variants [48, 49, 50]. Although it is less suited for direct integration, the Baikov representation has the advantage that implementing cuts is straightforward [48, 51, 52], and it is more compatible with twisted cohomology theory [53, 54, 55, 56]. The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc.
While it is evident that these two representations should be equivalent, since both are derived from the momentum representation, it was not clear how to directly derive one representation from the other. This problem is addressed in this letter. It is demonstrated that these two representations are related by a simple transformation.
This letter is organized as follows. A transformation on homogeneous polynomials is introduced in section II. The transformation is then applied to parametric Feynman integrals in section III. It is proven that the two representations related through this transformation are dual to each other. The equivalence between the dual representation and the Baikov representation is proven in section IV.
@@ II Polynomial transformation
We assume that the degree of $\mathcal{F}$ is $L+1$. Then by virtue of the homogeneity of $\mathcal{F}$, we have
@@
Taking the derivatives of both sides with respect to $u_{i}$, we get
@@
Generally speaking, the transformation defined by eq. II.1a may not be reversible. Even if it is, the obtained function $\mathcal{G}(u)$ may not be a rational function. However, for the $\mathcal{F}$ polynomial of a loop integral, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
@@ III Representation transformation
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\\ &\times\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
@@
$w_{\lambda}(x)\equiv e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v+i0^{+})^{\lambda+1}}e^{-ivx}~{}.$ (III.2)
@@
$w_{-1}(x)=$ $~{}2\pi\delta(x)~{},$
$w_{\lambda}(x)=$ $~{}2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
The Fourier transform of $w_{\lambda}(x)$ gives
$\frac{1}{x^{\lambda+1}}=e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(v)e^{-ivx},~{}\mathrm{Im}\{x\}<0~{}.$ (III.3)
Obviously we can write $\frac{1}{\Gamma(\lambda+1)}\int_{0}^{\infty}\mathrm{d}xx^{\lambda}$ as $\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}xw_{\lambda}(x)$. By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$] by using eqs. (III.2 and III.3), the integral $I_{x}$ in eq. (III.1) becomes
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(-2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\\ &\times\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(-2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\\ &\times\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
where $\mathrm{d}\Pi_{v}^{(n+1)}\equiv\prod_{i=1}^{n+1}\mathrm{d}v_{i}~{}\delta(1-\mathcal{E}^{(1)}(v))$. The last step of this equation can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
@@
$\Lambda_{i,j(n+1)}=$ $~{}0,~{}j\neq n+1~{},$ (III.6a)
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $~{}n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
with $n_{0}$ a constant. As can be seen in appendix B, a family of loop integrals with a complete set of propagators do satisfy these conditions. That is, the completeness of propagators implies the completeness of the corresponding $\mathcal{F}$ polynomial. The first condition of eqs. (III.6) is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$. For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eqs. (III.6), it can be proven that (see appendix A)
$\begin{split}&\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&~{}(-2i\pi)^{n+1}e^{\frac{i}{2}\pi(L+1)\lambda_{0}}C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{\lambda_{0}^{\prime}}~{},\end{split}$ (III.7)
where $\lambda_{0}^{\prime}\equiv-\lambda_{0}-\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a ($v$-independent) constant. Substituting eq. (III.7) into eq. (III.4) gives
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&\exp\left(\frac{i\pi}{2}\left[(L+1)\lambda_{0}+\sum_{i=1}^{n+1}\lambda_{i}\right]\right)\\ &\times C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(-\lambda_{0}^{\prime})}~{}I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$. The integral $I_{u}$ is defined by
$I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ (III.9)
$\equiv$ $~{}\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}~{}.$
@@
$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)=e^{\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{-iux}$
Table 1: The duality between the two representations. Here the $\frac{1}{u^{\lambda+1}}$ should be understood as $\frac{1}{(u+i0^{+})^{\lambda+1}}$.
@@ IV The Baikov representation
We consider a loop integral with propagators
@@
where $l$ are the loop momenta and $p$ are the external momenta. For simplicity, we assume that the Gram determinant of the external momenta $p$ does not vanish, and all the propagators are complete and linearly independent. Here, by complete, we mean that $v_{i}$ form a complete basis such that all the scalar products $l_{i}\cdot l_{j}$ and $l_{i}\cdot p_{j}$ can be expressed in terms of $v_{i}$. For future convenience, we collectively denote $l$ and $p$ by $q$. Scalar integrals can be parametrized by integrals of the form [44, 45]
$\begin{split}&J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\\ &\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},\end{split}$ (IV.2)
where $E$ is the number of the external momenta $p$. Here we have omitted some irrelevant prefactors for brevity. Notice that we use such a convention that the indices of the propagators are $-\lambda_{i}-1$ rather than the more conventional $-\lambda_{i}$. The polynomial $P(v)$ is proportional to the Gram determinant of the momenta $q$. That is,
@@
In this section, we will prove that $J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})$ is equivalent to the integral $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ in eq. III.9 up to some constants.
Taking $\mathcal{E}^{(1)}(u)=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
@@
$=$ $\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n+1}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\Gamma(-\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
×$\times\left[\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n+1}^{\prime}+1)}+\mathcal{G}_{-}^{\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n+1}^{\prime}+1)}\right]$
$=$ $2\Gamma(-\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\frac{-4i\pi}{\Gamma(1+\lambda_{0}^{\prime})}\cos\left(\pi\lambda_{n+1}^{\prime}\right)$
×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
where $\mathcal{G}_{+}\equiv\left.\mathcal{G}\right|_{u_{n+1}=1}$ and $\mathcal{G}_{-}\equiv\left.\mathcal{G}\right|_{u_{n+1}=-1}$. Due to the positivity of $|u_{n+1}|$, the integral receives contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$. The second step of this equation is carried out by replacing the integration variables with $x_{i}\to-x_{i}$ for the first term in the square bracket. In the last step, we have closed the contour around the branch cut. As a result, only the discontinuity of $\mathcal{G}_{-}^{\lambda_{0}^{\prime}}$ contributes to the integral. The discontinuity arises from the region where $\mathcal{G}_{-}<0$, which gives rise to the theta function $\theta\left(-\mathcal{G}_{-}\right)$. Notice that we use such a convention that $\mathrm{Im}\{\mathcal{F}\}<0$. Thus $\mathrm{Im}\{\mathcal{G}\}>0$. In the last step we have expressed $\sin(\pi\lambda_{0}^{\prime})$ in terms of $\frac{-\pi}{\Gamma(-\lambda_{0}^{\prime})\Gamma(1+\lambda_{0}^{\prime})}$. Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=-\lambda_{0}-\frac{1}{2}(L+E+1)$. Comparing the above integral with the Baikov integral in eq. IV.2, we see that these two integrals are equivalent if
$\begin{split}P(v_{1},v_{2},\dots,v_{n})=&-\mathcal{G}_{-}\\ =&-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.\end{split}$ (IV.4)
@@
$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6a)
$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.6b)
$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.6c)
$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.6d)
@@
$\begin{split}\mathcal{G}=~{}&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.7)
@@
$P=~{}$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
[…]
Solving eq. (IV.1) for $l_{i}\cdot l_{j}$ and $l_{i}\cdot p_{j}$, we get
@@
$\begin{split}P=&\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})\right.\\ &\left.-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.\end{split}$ (IV.8)
Comparing eq. IV.7 with eq. IV.8, we get eq. IV.4.